OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-5,-4) and (7,-5).

OpenStudy (anonymous):

@Hero

OpenStudy (aum):

First find the slope of the line from the two given points. If \((x_1,y_1)\) and \((x_2,y_2)\) are the two given points, then the slope \(\Large m = \frac{y_2-y_1}{x_2-x_1}\). What is the slope m?

OpenStudy (texaschic101):

you figure out the slope yet ?

OpenStudy (masumanwar):

We can solve this equation for y= mx+c

OpenStudy (anonymous):

@aum @texaschic101 @masumanwar does m=5

OpenStudy (aum):

(-5,-4) and (7,-5) are the two given points. \(\Large m = \frac{y_2-y_1}{x_2-x_1} = \frac{-5-(-4)}{7-(-5)} = \frac{-5+4}{7+5}=?\)

OpenStudy (anonymous):

that equals m=-1/12 @aum im sorry i was busy chattin with my fans

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

can u help meh pls *-*

OpenStudy (anonymous):

well @aum does the computation for you, all you have to do is perform the arithmetic operations to solve for m, then using any of the points given... using the point-slope formula of a line\[(y-y_1)=m(x-x_1)\]you can have the equation of the line...

OpenStudy (anonymous):

wut im xD

OpenStudy (anonymous):

can u explain that 2 meh @Orion1213

OpenStudy (anonymous):

the two given points as we all know form the line, both of them lie on the line that we are looking for so they both satisfy the equation of the line... there is only one slope m, that you have obtained already... plug-in m and the coordinates of the point in place of \(x_1\) and \(y_1\) in the point-slope formula i gave above...

OpenStudy (anonymous):

so is it y=1/12x-or +b

OpenStudy (anonymous):

your slope is negative i think...

OpenStudy (anonymous):

oh ok so is it y=1/12x-b

OpenStudy (anonymous):

the final form is the slope-intercept form wherein you are familiar\[y=mx+b\]

OpenStudy (anonymous):

let us check: \(P_1~(-5,-4)\) and \(m=-1/12\)\[(y-(-4))=-\frac{1}{12}(x-(-5))\]\[y+4=\frac{-x-5}{12}\]\[y=-\frac{1}{12}x-4\frac{5}{12}\]\[y=-\frac{1}{12}x-\frac{53}{12}\]

OpenStudy (anonymous):

you must get the same equation once you plug-in the other point (7,-5)

OpenStudy (anonymous):

hmmm interesting soo sooo y=-1/12x-53/12 is tha answer or do I have 2 divide

OpenStudy (anonymous):

Write the equation 5x - y + 6 = 0 in standard form with integer coefficients. @Orion1213 can u help meh with this pls

OpenStudy (anonymous):

is it Y=5X+6

OpenStudy (anonymous):

Ax + By + C = 0 is the standard form of a line...

OpenStudy (anonymous):

A, B and C are the integer coefficients...